OpenStudy (anonymous):

I will fan and medal!!! Just guide me through the steps please :) The length of a diagonal of a square sandbox is 9 feet. What is the approximate length of one side of the sandbox?

OpenStudy (anonymous):

ok so you have a square box and the diagonal is 9 right?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Do you know what the length of the entire outside is?

OpenStudy (anonymous):

ok and diagonal is the samelength of a side right?

OpenStudy (anonymous):

The diagonal is not the same as one side.

OpenStudy (anonymous):

No @ScarlettBlack

OpenStudy (anonymous):

no its square root of (10 spuared +10 squared )

OpenStudy (anonymous):

so the side length is 20

OpenStudy (anonymous):

@wakeboarder Thats not one of my answer choices tho

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

it should be around 6 or 6.3

OpenStudy (anonymous):

no it should be 14

OpenStudy (anonymous):

d=square root of 2 * a Solving for a a=2 d/2 = square root of 2 * 9/2 = 6.36

OpenStudy (anonymous):

Can I see your choices?

OpenStudy (anonymous):

oh i was using 10

OpenStudy (anonymous):

yeah its 6.36

OpenStudy (anonymous):

Unfortunately I have to go now.

OpenStudy (anonymous):

no its wrong

OpenStudy (anonymous):

about 12.5

OpenStudy (anonymous):

ok sorry. :) Here are my choices : A. 4.5 ft B. 5.2 ft C. 6.4 ft D. 12.7 ft

OpenStudy (anonymous):

d.12.7